OpenStudy (anonymous):

The perimeter of the rectangle is 8x -4. The length of side YZ is 2x – 3. What is the length of side XY?

OpenStudy (photon336):

the sum of all the sides must equal 8x-4

OpenStudy (photon336):



OpenStudy (anonymous):

would it combine to 4x-6? @Photon336

OpenStudy (photon336):

sorry i made a mistake earlier (8x-4)-(4x-6) = 4x+2 (4x+2)+(4x-6) = 8x-4 the value of the other side should be 2x+1 (4x+2)*(1/2) = 2x+1

OpenStudy (photon336):

because (4x-6) + (4x+2) = 4x+4x = 8x -6+2 = -4 so 8x-4 so the value of the side you're looking for should be 2x+1

OpenStudy (photon336):

@frebrezze

OpenStudy (photon336):

the perimeter is the sum of all the sides, so the sum is 8x-4 so you need to subtract 2x-6 from that to find the value of the other two sides.

OpenStudy (photon336):

@frebrezze does this make sense?